# How connection probability shapes fluctuations of neural population dynamics

### v1

@@ Appendix A Mesoscopic dynamics in the presence of heterogeneous external drive
The random variables $\hat{\mu}_{i}$ for each neuron $i$ have mean zero and are independently and identically distributed with probability density $\rho(\hat{\mu})$. They can be interpreted as heterogeneity of the external drives, the resting potentials or the thresholds $\vartheta$ of the different neurons. Importantly, the variables $\hat{\mu}_{i}$ are constant in time, and thus represent quenched disorder. The derivation of the microscopic annealed network (12) does not explicitly depend on the value of $\mu(t)$ and we therefore obtain the same results with $\mu(t)$ replaced with Eq. (49). Specifically, Eq. (15) is modified as follows
@@
[…]
@@
Appendix B Estimation of the rate variance from population activities
The variance of the population firing rate is not directly accessible from measurements in real biological networks. However, we can esimate the rate variance by the variance of the emphirical population activity. We assume that in small time bins $(t,t+\Delta t]$, neurons fire spikes independently with conditional intensities $\lambda_{i}(t)$ given the past. Then, from the law of total variance and the conditionally Poisson statistics, we have for the variance of the total number of spikes in a time bin
[…]
Appendix C Condition for the validity of the firing-rate approximation (26)
@@
In Eq. (26), a simple approximation for the fixed-point firing rate is given, namely $r_{0}\approx-\mu_{0}/w$, which we used for an inhibitory network ($w<0$) in the non-saturated regime ($0<\mu_{0}\leq-wr_{\text{m}}$). This approximation corresponds to the solution of the 1st-order MF model in the limit $\beta\to\infty$. We argued that under certain sufficient conditions, the approximation is good for the 2nd-order MF model as well. Here we provide a justification for these conditions. For biological reasons, we make a slightly stronger assumption on the working point of our system: we assume that the fixed point is in the convex part of the transfer function, i.e. $h_{0}<0$ and $r_{0}<r_{\text{m}}/2$, as drawn in Fig. 3.
@@
For the firing-rate approximation to be applicable also to the 2nd-order MF model, we require that the relative difference between the firing rates of the 1st- and 2nd-order MF theory is small:
@@
[…]
@@ Appendix D Partial derivatives of the function $F$ and approximations thereof
The approximate expressions Eq. (28) and Eq. (27) for $h$ and $\sigma^{2}$ at the fixed point can be used for further simplifications of the partial derivatives:
@@ Appendix E Linear stability analysis
[…]
@@
[…]
Both must be satisfied simultaneously. Here, atan2 is the two-argument arctangent and $k\in\mathbb{N}_{0}$. In the $(d,w)$-parameter space, these conditions correspond to a family of curves, one curve for each $k$. At the $k$-th curve an oscillatory perturbation with frequency $\omega_{k}$ becomes unstable. For the overall instability boundary it is sufficient that at least one mode becomes unstable. We found empirically that the overall instability boundary is given by the mode $k=0$, whereas higher modes yield regions of instability that are contained already in the instability region given by $k=0$. For the solution of both conditions a numerical solution of the fixed point equations is necessary. In Fig. 4b) we scan through each parameter point and solve the phase condition for $\omega$. In the investigated parameter range we were always able to find a unique solution for the phase condition for each mode. We then determined the boundary in the $(d,w)$-space that separated the regions where the right-hand side of the amplitude condition, Eq. (56), is larger and smaller than $1$, respectively. The equations are more simple for the 1st-oder MF model where $\hat{F}_{\sigma}=0$. The simplified amplitude condition reads in this case
@@ Appendix F Susceptibility matrix
$\tilde{\operatorname{\chi}}_{11}$ $=\frac{1}{\mathcal{D}}\left(i\omega+\frac{2}{\tau}-F_{\sigma}\hat{\beta}e^{-i\omega d}\right)$ (57a)
$\tilde{\operatorname{\chi}}_{12}$ $=\frac{1}{\mathcal{D}}F_{\sigma}\hat{\alpha}e^{-i\omega d}$ (57b)
$\tilde{\operatorname{\chi}}_{13}$ $=\frac{1}{\mathcal{D}}\hat{\gamma}\hat{\alpha}e^{-i\omega d}\frac{i\omega+\frac{2}{\tau}}{i\omega+\frac{1}{\tau}}$ (57c)
$\tilde{\operatorname{\chi}}_{21}$ $=\frac{1}{\mathcal{D}}F_{h}\hat{\beta}e^{-i\omega d}$ (57d)
$\tilde{\operatorname{\chi}}_{22}$ $=\frac{1}{\mathcal{D}}\left(i\omega+\frac{1}{\tau}-F_{h}\hat{\alpha}e^{-i\omega d}\right)$ (57e)
$\tilde{\operatorname{\chi}}_{23}$ $=\frac{1}{\mathcal{D}}\hat{\gamma}\hat{\beta}e^{-i\omega d}$ (57f)
@@
$\tilde{\operatorname{\chi}}_{33}$ $=\frac{1}{\mathcal{D}}\left(i\omega+\frac{2}{\tau}\right)-\frac{1}{\mathcal{D}}F_{\sigma}\hat{\beta}e^{-i\omega d}$ (57h)
$-\frac{1}{\mathcal{D}}\frac{i\omega+\frac{2}{\tau}}{i\omega+\frac{1}{\tau}}F_{h}\hat{\alpha}e^{-i\omega d}$ (57i)
@@
$\mathcal{D}(\omega)$ $=\left(i\omega+\frac{1}{\tau}\right)\left(i\omega+\frac{2}{\tau}\right)-\left(i\omega+\frac{2}{\tau}\right)F_{h}\hat{\alpha}e^{-i\omega d}$
$\qquad-\left(i\omega+\frac{1}{\tau}\right)F_{\sigma}\hat{\beta}e^{-i\omega d}$
$\hat{\alpha}$ $=\frac{w}{\tau},\quad\hat{\beta}=\frac{w^{2}}{\tau^{2}}\left(\frac{1}{C}-\frac{1}{N}\right),\quad\hat{\gamma}=\frac{1}{\sqrt{N}}.$

### v2

@@ VI Acknowledgment
VII Data availability
The code that supports the findings of this article is openly available [87].
@@ Appendix A Mesoscopic dynamics in the presence of heterogeneous external drive
[…]
@@
$\begin{aligned} \tau\frac{\mathrm{d}\bar{x}}{\mathrm{d}t}&=-\bar{x}+\mu(t)+w\left[r(t-d)+\sqrt{\frac{r(t-d)}{N}}\eta(t)\right]\\ \tau\frac{\mathrm{d}\sigma_{x}^{2}}{\mathrm{d}t}&=-2\sigma_{x}^{2}+\frac{w^{2}\left(1-p\right)}{\tau pN}r(t-d)\\ \tau\frac{\mathrm{d}\xi}{\mathrm{d}t}&=-\xi+\sqrt{2\tau G\left(\bar{x}(t),\sigma^{2}(t)\right)}\zeta(t)\end{aligned}$
@@
Appendix B Fixed point solutions
At equilibrium, the sparse limit, Eq. (22), gives
$\bar{h}_{0}=\mu_{0}+wr_{0},\quad\sigma_{0}^{2}=\frac{w^{2}}{2\tau C}r_{0},\quad r_{0}=F(\bar{h}_{0},\sigma_{0}^{2})$ (51)
which yields the fixed-point equation (25) for $h_{0}$.
Figure 8: Mean stationary population rate of the quenched and annealed networks and fixed-point solutions of the noiseless mean-field dynamics. Population averaged firing rates as a function of (a) the external input $\mu_{0}$, (b) the connection probability $p$ for $C=100$ constant, and (c) the coupling strength $w$. Orange (blue) line: MF1 (MF2), brown line: MFsparselim; red squares (green circles): annealed (quenched) network simulation; dotted line: analytical approximation (26). The inset in (c) shows the same data, but for a larger range of $w$ in a double logarithmic scaling. Gray dashed line: 2nd-order theory $r_{\infty}$ for the firing rate in the limit $w\to$-$\infty$ (Eq. (55)). Note: MFsparselim is identical to MF2 in (a) and (c). Parameters: $N=1000$, $\beta=5$ mV-1 (a) $p=0.1$ and w=1w=-1 mVs, (b) $\mu_{0}=10$ mV, w=1w=-1 mVs, (c) $p=0.1$, $\mu_{0}=10$ mVs.
Because we assume an inhibitory network ($w<0$), the linear function on the left-hand side of Eq. (25) has a negative slope. If we choose $w$ such that the intersection occurs for values $\bar{h}_{0}$ lower than the inflection point $\vartheta=0$ of the hazard function $\phi(h)$ (corresponding to the biologically relevant convex part of $\phi(h)$), then from the graphical solution it is clear that the 2nd-order MF dynamics exhibits a fixed-point with lower mean input potential $\bar{h}_{0}$ and higher population firing rate $r_{0}$ compared to the 1st-order MF dynamics. This effect becomes stronger for smaller connection probability $p$.
The mean stationary firing rates of the quenched and annealed microscopic models are well captured by the numerical solution of the fixed-point equation (25) (Fig. 8a-c). As a function of the external drive $\mu_{0}$, we observe a strikingly linear dependence between the saturation regimes at low and high values of $\mu_{0}$ (Fig 8a). Furthermore, we find that the firing rate is roughly constant when the connection probability $p$ is varied while $N$ is fixed (Fig 8b). The weak dependence on $p$ is slightly better predicted by the 2nd-order MF theory than by the 1st-order MF theory, which, by definition, has no dependence on $p$. We also observe a power-law behavior of the mean firing rate in both microscopic simulations and fixed-point solutions if the coupling strength is not too strong (Fig. 8c). We note that, in general, the relative error of the 1st-order MF dynamics (and hence the mean-connectivity network) with respect to the stationary mean population rate is rather small compared to the much larger error with respect to the second-order statistics reported in Fig. 1.
B.1 Analytical approximation of fixed points
The linear dependence on $\mu_{0}$, the weak dependence on $p$ and the power-law dependence on $w$ can be explained with our MF theory as follows. Firstly, we notice that for a sufficiently large steepness $\beta\gg 1/\sigma_{0}^{2}$ of the single-neuron transfer function $\phi$, the 2nd-order MF dynamics remains practically unaltered if we take the limit $\beta\to\infty$. In this limit, the single-neuron transfer function tends to $r_{\text{m}}\theta(\bar{h})$, where $\theta$ denotes the Heaviside step function. Indeed, for the values of $\beta$ considered in this paper, taking the limit $\beta\to\infty$ does not noticeably alter the effective nonlinearity $F$ and hence the 2nd-order MF dynamics. The limiting value of the fixed point can be calculated analytically from the graphical representation (Fig. 3a) which yields an approximate analytical expression for the firing rate in the 1st-order MF dynamics,
This analytical expression predicts that $r_{0}$ is inversely proportional to the coupling strength in the balanced regime $0<\mu_{0}<-wr_{\text{m}}$. Indeed, the numerical fixed-point solution for the 1st-order MF dynamics exhibits an excellent quantitative agreement with this prediction (Fig. 8c).
Secondly, we note from the geometry of the graphical solution of the fixed points that the intersection point of the 1st-order MF dynamics serves as a rough approximation for the firing rate $r_{0}$ of the 2nd-order MF dynamics as well (Fig. 3a) unless $w$ is extremely negative as discussed below. Indeed, the horizontal stretching of the transfer function in the 2nd-order theory mainly affects the value of the fixed point $\bar{h}_{0}$ rather than the firing rate $r_{0}$. More precisely, a sufficient condition for the validity of our approximation is that $\mu_{0}/w\gg(1-p)/(\tau C)$ and $\mu_{0}/w>0.023r_{\text{m}}$ (see Appendix B.3). We expect that the approximation (26) slightly underestimates the firing rates if the intersection is below the inflection point ($\bar{h}_{0}<0$) and slightly overestimates the firing rates if the intersection is above the inflection point ($\bar{h}_{0}>0$). Our approximation Eq. (26) well explains the piecewise linear behavior of the firing rate as a function of the external input as observed in Fig. 8a. Furthermore, Eq. (26) also explains why the firing rates are roughly independent of the connection probability $p$ (Fig. 8b), and why we see the power-law $w^{-1}$ even in non-fully connected networks (Fig. 8c) unless $w$ is extremely negative (this case is treated below). In general we note that, compared to the 1st-order MF dynamics, the 2nd-order MF dynamics shows a better agreement with the firing rate in microscopic simulations, in particular, it fits the annealed network almost perfectly.
$\sigma_{0}^{2}\approx-\frac{w\mu_{0}(1-p)}{2\tau Np}$ (53)
B.2 Limit of strongly inhibitory networks
The graphical solution of this equation clearly shows a vanishing firing rate in the 1st-order MF theory ($p\to 1$), whereas the limiting equation has a non-vanishing solution for $p<1$ (2nd-order MF theory, Fig. 8c, inset). The non-vanishing, limiting firing rate in the 2nd-order MF theory is in good agreement with simulations of the annealed and quenched microscopic model. In contrast, the 1st-order MF theory generally underestimates the firing rates in strongly inhibitory networks: for $w\to$-$\infty$, the 1st-order MF dynamics has a rate that converges to zero like $\mu_{0}/w$ as predicted by Eq. (26).
B.3 Condition for the validity of the firing-rate approximation (26)
@@
In Eq. (26), a simple approximation for the fixed-point firing rate is given, namely $r_{0}\approx-\mu_{0}/w$, which we used for an inhibitory network ($w<0$) in the non-saturated regime ($0<\mu_{0}\leq-wr_{\text{m}}$). This approximation corresponds to the solution of the 1st-order MF theory in the limit $\beta\to\infty$. We argued that under certain sufficient conditions, the approximation is good for the 2nd-order MF dynamics as well. Here we provide a justification for these conditions. For biological reasons, we make a slightly stronger assumption on the working point of our system: we assume that the fixed point is in the convex part of the transfer function, i.e. $h_{0}<0$ and $r_{0}<r_{\text{m}}/2$, as drawn in Fig. 3a.
@@
For the firing-rate approximation to be applicable also to the 2nd-order MF dynamics, we require that the relative difference between the firing rates of the 1st- and 2nd-order MF theory is small:
@@
Here, $r_{0}^{\text{1st}}=-\mu_{0}/w$ and $r_{0}^{\text{2nd}}=F(h_{0},\sigma_{0}^{2})$. Furthermore, we have used that $r_{0}^{\text{2nd}}>r_{0}^{\text{1st}}$ because of our assumption, $h_{0}<0$ (Fig. 3a). To estimate the left hand side of (56) from above, we want to obtain an upper bound approximation of the rate $r_{0}^{\text{2nd}}=(h_{0}-\mu_{0})/w$. To this end, we lower bound the input potential $h_{0}$ (note that $w<0$). Because $h_{0}$ is expected to be on the order of the width $\sigma_{0}$ of the transfer function, we write the lower bound as $a\sigma_{0}<h_{0}$, where $a$ is a positive number that is still unknown. This lower bound of $h_{0}$ yields an upper bound estimate for the mean firing rate: $r_{0}^{\text{2nd}}<(-a\sigma_{0}-\mu_{0})/w$. Hence, the condition (56) is surely met if
@@
Appendix C Estimation of the rate variance from population activities
The variance of the population firing rate is not directly accessible from measurements in real biological networks. However, we can estimate the rate variance by the variance of the empirical population activity. We assume that in small time bins $(t,t+\Delta t]$, neurons fire spikes independently with conditional intensities $\lambda_{i}(t)$ given the past. Then, from the law of total variance and the conditionally Poisson statistics, we have for the variance of the total number of spikes in a time bin
where $r(t)=N^{-1}\sum_{i=1}^{N}\lambda_{i}(t)$ is the population firing rate and $\left\langle\cdot\right\rangle$ denotes the trial average. Hence,
@@ Appendix D Partial derivatives of the function $F$ and approximations thereof
The approximate expressions Eq. (27) and Eq. (28) for $h_{0}$ and $\sigma_{0}^{2}$ at the fixed point can be used for further simplifications of the partial derivatives:
@@ Appendix E Linear stability analysis
$\begin{aligned} \delta h(t)=\hat{h}(\lambda)e^{\lambda t}\\ \delta\sigma^{2}(t)=\hat{\sigma}(\lambda)e^{\lambda t},\end{aligned}$
@@
$-\text{atan2}\left(2-\hat{F}_{\sigma}\cos(\omega d),\omega\tau+\hat{F}_{\sigma}\sin(\omega d)\right).$
Both must be satisfied simultaneously. Here, atan2 is the two-argument arctangent and $k\in\mathbb{N}_{0}$. In the $(d,w)$-parameter space, these conditions correspond to a family of curves, one curve for each $k$. At the $k$-th curve an oscillatory perturbation with frequency $\omega_{k}$ becomes unstable. For the overall instability boundary it is sufficient that at least one mode becomes unstable. We found empirically that the overall instability boundary is given by the mode $k=0$, whereas higher modes yield regions of instability that are contained already in the instability region given by $k=0$. For the solution of both conditions a numerical solution of the fixed point equations is necessary. In Fig. 3b we scan through each parameter point and solve the phase condition for $\omega$. In the investigated parameter range we were always able to find a unique solution for the phase condition for each mode. We then determined the boundary in the $(d,w)$-space that separated the regions where the right-hand side of the amplitude condition, Eq. (62), is larger and smaller than $1$, respectively. The equations are simpler for the 1st-order MF dynamics where $\hat{F}_{\sigma}=0$. The simplified amplitude condition reads in this case
@@ Appendix F Susceptibility matrix
$\tilde{\operatorname{\chi}}_{11}$ $=\frac{1}{\mathcal{D}}\left(i\omega+\frac{2}{\tau}-F_{\sigma}\tilde{\beta}e^{-i\omega d}\right)$ (63a)
$\tilde{\operatorname{\chi}}_{12}$ $=\frac{1}{\mathcal{D}}F_{\sigma}\tilde{\alpha}e^{-i\omega d}$ (63b)
$\tilde{\operatorname{\chi}}_{13}$ $=\frac{1}{\mathcal{D}}\tilde{\gamma}\tilde{\alpha}e^{-i\omega d}\frac{i\omega+\frac{2}{\tau}}{i\omega+\frac{1}{\tau}}$ (63c)
$\tilde{\operatorname{\chi}}_{21}$ $=\frac{1}{\mathcal{D}}F_{h}\tilde{\beta}e^{-i\omega d}$ (63d)
$\tilde{\operatorname{\chi}}_{22}$ $=\frac{1}{\mathcal{D}}\left(i\omega+\frac{1}{\tau}-F_{h}\tilde{\alpha}e^{-i\omega d}\right)$ (63e)
$\tilde{\operatorname{\chi}}_{23}$ $=\frac{1}{\mathcal{D}}\tilde{\gamma}\tilde{\beta}e^{-i\omega d}$ (63f)
@@
$\tilde{\operatorname{\chi}}_{33}$ $=\frac{1}{\mathcal{D}}\left(i\omega+\frac{2}{\tau}\right)-\frac{1}{\mathcal{D}}F_{\sigma}\tilde{\beta}e^{-i\omega d}$ (63h)
$-\frac{1}{\mathcal{D}}\frac{i\omega+\frac{2}{\tau}}{i\omega+\frac{1}{\tau}}F_{h}\tilde{\alpha}e^{-i\omega d}$ (63i)
@@
$\mathcal{D}(\omega)$ $=\left(i\omega+\frac{1}{\tau}\right)\left(i\omega+\frac{2}{\tau}\right)-\left(i\omega+\frac{2}{\tau}\right)F_{h}\tilde{\alpha}e^{-i\omega d}$
$\qquad-\left(i\omega+\frac{1}{\tau}\right)F_{\sigma}\tilde{\beta}e^{-i\omega d}$
$\tilde{\alpha}$ $=\frac{w}{\tau},\quad\tilde{\beta}=\frac{w^{2}}{\tau^{2}}\left(\frac{1}{C}-\frac{1}{N}\right),\quad\tilde{\gamma}=\frac{1}{\sqrt{N}}.$
@@ Appendix I There exists exactly one fixed point for inhibitory networks with positive external drive
Appendix J Summary tables
An overview of the parameters and the population models in different parameter regimes are provided in Table  2 and Table 3, respectively.
Parameter Value / Relations Description
$\tau$ 2020 ms time constant of the low-pass filter dynamics
$r_{\text{m}}$ 100100 Hz maximal firing rate
$\beta$ 55 mV-1 steepness of the single-neuron transfer function (threshold sharpness)
$\vartheta$ 0 mV inflection point of single-neuron transfer function (soft threshold)
$d$ 0 ms (unless otherwise noted) transmission delay
$p$ see figure captions connection probability
NN see figure captions number of neurons
$C$ $C=pN$ number of incoming synapses
[…]
$J$ $J=w/C$, (mVs) efficacy of single synapse
$a_{ij}$ $\left(a_{ij}\right)\sim\text{Uniform}\left(\left\{\mathcal{A}\in\{0,1\}^{N\times N}:\sum_{j=1}^{N}\mathcal{A}_{ij}=C\right\}\right)$ adjencency matrix, $a_{ij}=\mathbbm{1}_{\{\text{neuron $j$ connected to $i$}\}}$
$J_{ij}$ $J_{ij}=Ja_{ij}$ synaptic efficacy from neuron $j$ to $i$
$w_{ij}$ $w_{ij}=NJ_{ij}$, $\left\langle w_{ij}\right\rangle=w$ rescaled synaptic efficacy
Table 2: Model parameters and their values.
Parameter regime MF1, Eq. (21) MF2, Eq. (20) MFsparselim, Eq. (22)
$N,C\to\infty$, $p=C/N=O(1)$ (dense) $\tau\frac{\mathrm{d}}{\mathrm{d}t}\bar{h}=-\bar{h}+\mu_{\rm ext}(t)+wr(t-d)\ ,\ \ r(t)=\phi\bigl(\bar{h}(t)\bigr)$ (74) $\bullet$ exact
$N\to\infty$, $C=O(1)$ (sparse) $\to$ Eq. (74)
✓ stationary population rate (for not too strong coupling)
✗ time-dependent rate response poor
$\tau\frac{\mathrm{d}}{\mathrm{d}t}\bar{h}$ $=-\bar{h}+\mu_{\rm ext}(t)+wr(t-d)$ (75) $\tau\frac{\mathrm{d}}{\mathrm{d}t}\sigma^{2}$ $=-2\sigma^{2}+\frac{w^{2}}{\tau C}r(t-d),$ (75) $r(t)$ $=F(\bar{h}(t),\sigma^{2}(t))$ (75) $\bullet$ exact for annealed network
$\bullet$ for quenched network:
✓ stationary population rate better than MF1
✓ dynamical rate response
$N<\infty$, $p=1$ $\tau\frac{\mathrm{d}\bar{h}}{\mathrm{d}t}$ $=-\bar{h}+\mu_{\rm ext}(t)+w\left[r(t-d)+\sqrt{\frac{r(t-d)}{N}}\eta(t)\right]$ (76) $r(t)$ $=\phi\bigl(\bar{h}(t)\bigr)$ (76) $\bullet$ exact $\to$ Eq. (75) with $C=N$
✗ no finite-size fluctuations
✗ stationary and time-dependent rate poor
$N<\infty$, $0<p<1$ $\to$ Eq. (76)
$\bullet$ does capture finite-size fluctuations but with significant errors for annealed and (much more) for quenched networks
✓ captures finite-size fluctuations very well for annealed network
✓ strong improvement for quenched network compared to MF1
$\to$ Eq. (75)
✗ no finite-size fluctuations
Table 3: Performance summary of the mean-field theories for different parameter regimes. Exact, apart from the minor diffusion approximation of the shot noise, which is valid for sufficiently large $C$ and which is actually not really necessary: one could go back and replace the term $A(t)\approx r(t)+\sqrt{r(t)/N}\eta(t)$ by a Poissonian shot noise $A(t)=\text{Pois}(Nr(t)dt)/(Ndt)$, which would recover the exact dynamics.